# Learning Autonomous Surgical Irrigation and Suction with the da Vinci Research Kit Using Reinforcement Learning

### v1

@@ VI-D Real-World Performance of Combined Irrigation-Suction
As discussed in Section V-C, during the complete autonomous irrigation-suction experiments, the weight of the contents inside the container is measured three times: the initial weight, the weight after autonomous irrigation and suction, and the weight after an additional manual suction for cleanup. The results are shown in Fig. 14. Snapshots from the experiments are shown in Fig. 17.
@@
We further calculate the difference between the weight after autonomous irrigation-suction and the one after manual cleanup, resulting in an average value of 2.42 (± $\pm\,2.04$). This value represents the amount of liquid that can be suctioned but is not actually removed during autonomous suction, reflecting the autonomous suction performance, which is similar to the results obtained from the suction-only trials. However, we observe more back-and-forth motion during suction this time, compared to the suction-only trials, likely due to the presence of the dark-colored ketchup that is not fully diluted.
@@ VI-E Suboptimal Outcomes
For suction, the typical failure case is that the agent fails to navigate the EE to a pool of liquid, which is more likely to occur when the pool is small. In the example shown in Fig. 18b, the EE was navigated to the top-left side initially but continued to move to the lower-right before removing all liquid in the top-left side. The agent eventually attempted to move the EE back to the top-left but failed due to the EE touching the container and violating the force constraint, as will be discussed in Section VII-C. In the trial shown in Fig. 18c, no attempt was even made to guide the EE to the remaining liquid pool. This is barely seen during the evaluation in the simulator, but more frequent in the real world, which can be related to the sim-to-real gap.
@@ VII-A Sim-to-Real Gap
Various factors can be responsible for the gap between the simulation and the real world, making the agents’ performance degrade when compared with the results in the simulator. Since the RGB image is used in this work, the difference between the real-world image and the rendered image in the simulator can be a major source of the sim-to-real gap. Fig. 19 shows a rendered image from the simulator and one from a real-world trial, in the actual observation size (84 $\times$ 84). Although the colors and the camera pose are randomized during training, there are uncaptured discrepancies between simulation and reality, such as the bright reflections.
The difference in physics can also enlarge the sim-to-real gap. For example, suction is simulated by a cone-shaped force field, whereas it is achieved primarily by pressure differences in reality. In practice, the EE has to be in contact with the liquid surface to suction it, while in the simulator particles can be pulled toward the EE even if there is a distance between them. In addition, a rigid body is used to simulate the tissue, whereas the tissue is emulated using playdough and is soft in reality. The choice of using a rigid body is related to the computational performance considerations, as will be discussed in the following section.
@@ VII-B Computational Performance Considerations
There are several considerations of computational performance. To allow efficient simulation, a soft body was not used to simulate the tissue although PhysX 5 allows soft body simulation using the finite element method (FEM). FEM is computationally intensive and requires more GPU resources and processing time. We could not create a large number of parallel training environments when using FEM soft bodies, nor were we able to achieve efficient training in terms of wall-clock time. Instead, a rigid body is used to simulate the tissue as a trade-off, although it may enlarge the sim-to-real gap.
We have also considered the observation of historical images to capture the temporal information. However, considering the large rollout buffer size and the usage of demonstration data, a much larger amount of memory would be needed. In the current setting, the memory usage is already more than 30 gigabytes with IL for irrigation, making it impractical to stack the image observations with historical ones. To solve this issue, either a much larger memory should be used, or ML-Agents must be modified to allow streaming of the rollout data.
@@ VII-C Other Limitations and Future Work
In this work, no depth information is provided to the agent. Ideally, it may be possible to achieve a better performance when depth is taken into account. However, it is not included in this work as it may increase the sim-to-real gap, due to the additional noises and inaccuracies of the real-world depth images. Future work may either include the depth observation directly or train the agent with a stereo camera that captures two images side-by-side.
One limitation of the real-world experiments is that a hardcoded EE force constraint is imposed. If the force limit is reached when the agent attempts to move the EE toward a position, such as when the EE touches the tissue, the trial is terminated for safety. Although the agent is penalized when the EE is in contact with the tissue during training, force limit violations are still seen during some suction trials. Future work may employ a better training strategy, such as rewarding the agent for lifting up the EE when it is in contact with the tissue.
Further extension of this work to actual surgeries will need additional development of realistic surgical scenes and re-training of the agents. Additional validations are required as well to assess the performance in realistic environments, such as using cadavers.
@@ VIII Conclusion
In this work, we examined the problem of autonomous irrigation and suction in MIS using RL. To allow sim-to-real transfer, a new surgical robot learning simulation framework was built using Unity and PhysX 5, which enables the development of the simulated learning environments for irrigation and suction. With DR and carefully designed reward functions, combined with CL and IL, two vision-based agents were trained to autonomously complete irrigation and suction. We found that the irrigation agent achieves a much better performance when a curriculum is used, and that IL helps improve the suction agent but not the irrigation one. We evaluated their performance in the real world and showed that they can achieve satisfactory results in most cases. Additionally, our simulation platform can be extended and used to build simulated training environments for other surgical tasks in the future.

### v2

@@ VI-D Real-World Performance of Combined Irrigation-Suction
As discussed in Section V-C, during the complete autonomous irrigation-suction experiments, the weight of the contents inside the container is measured three times: the initial weight, the weight after autonomous irrigation and suction, and the weight after an additional manual suction for cleanup, as shown in Fig. 12a. The results are shown in Fig. 12b. Snapshots from the experiments are shown in Fig. 15.
@@
We further calculate the difference between the weight after autonomous irrigation-suction and the one after manual cleanup, resulting in an average value of 2.42 (± $\pm\,2.04$). This value represents the amount of liquid that can be suctioned but is not actually removed during autonomous suction, reflecting the autonomous suction performance, which is similar to the results obtained from the suction-only trials. An ANOVA test does not find a significant statistical difference between the three groups (combined suction-irrigation, suction-only with around 20 grams of initial weight, and suction-only with around 30 grams of initial weight), with a $p$-value of 0.89. Since the $p$-value is high, there is insufficient evidence to suggest that suction performance differs across these scenarios, implying that it may be relatively stable. There is also no strong statistical evidence of correlation between the irrigation and the suction performances, based on the Pearson correlation coefficient and Spearman’s $\rho$ ($-0.2953$ and $-0.3939$) and their $p$-values ($0.4075$ and $0.2629$), between (1) the final weight after manual cleanup and (2) the weight difference before and after manual cleanup. However, we observe more back-and-forth motion during suction this time, compared to the suction-only trials, likely due to the presence of the dark-colored ketchup that is not fully diluted.
@@ VI-E Suboptimal Outcomes
For suction, the typical failure case is that the agent fails to navigate the EE to a pool of liquid, which is more likely to occur when the pool is small. In the example shown in Fig. 16b, the EE was navigated to the top-left side initially but continued to move to the lower-right before removing all liquid on the top-left side. The agent eventually attempted to move the EE back to the top-left but failed due to the EE touching the container and violating the force constraint, as will be discussed in Section VII-C. In the trial shown in Fig. 16c, no attempt was even made to guide the EE to the remaining liquid pool. This is barely seen during the evaluation in the simulator, but is more frequent in the real world. One possible reason is the sim-to-real gap, as discussed in Section VII-A.
@@ VII-A Sim-to-Real Gap
There are various discrepancies between the real-world environment and the simulated one. Since the RGB image is used in this work, one major discrepancy is the difference between the real-world image and the rendered image in the simulator. Fig. 17 shows a rendered image from the simulator and one from a real-world trial, in the actual observation size (84 $\times$ 84). Although the colors and the camera pose are randomized during training, there are uncaptured discrepancies between simulation and reality, such as bright reflections. Additionally, the liquid color inside the tissue may not be uniform after irrigation, while our suction training environment starts with uniform liquid color in each episode due to implementation simplicity. The simulated fluid color mixing effect may also not fully emulate that in the real world. These factors can cause agents to perform worse in the real world compared to their simulation performance. Certain factors, like reflections, could be better simulated and considered during training to obtain agents more capable of handling or inferring such information. Otherwise, techniques such as image-to-image translation and domain adaptation (DA) could be explored in future training.
The difference in physics can also enlarge the sim-to-real gap. For example, suction is simulated by a cone-shaped force field, whereas it is achieved primarily by pressure differences in reality. In practice, the EE has to be in contact with the liquid surface to suction it, while in the simulator, particles can be pulled toward the EE even if there is a distance between them. In addition, a rigid body is used to simulate the tissue, whereas the tissue is emulated using playdough and is soft in reality. The choice of using a rigid body is related to the computational performance considerations, as will be discussed in the following section.
We did not perform a quantitative analysis of sim-to-real transfer performance, including experiments on how well the model generalizes to different camera poses, lighting conditions, and object colors in the real world. Additionally, control frequency and latency, while not strictly part of the sim-to-real gap, can further impact real-world performance. To better understand the sim-to-real gap, future work could include further experiments and analysis comparing real-world performance under various conditions with that observed in the simulator.
@@ VII-B Computational Performance Considerations
There are several considerations of computational performance. To allow efficient simulation, a soft body was not used to simulate the tissue although PhysX 5 allows soft body simulation using the finite element method (FEM). FEM is computationally intensive and requires more GPU resources and processing time. We could not create a large number of parallel training environments when using FEM soft bodies, nor were we able to achieve efficient training in terms of wall-clock time. Instead, a rigid body is used to simulate the tissue as a trade-off. However, this simplification also results in the exclusion of key tissue characteristics such as viscoelastic behavior, especially when the tool is in contact with the tissue. Liquid permeability, absorption, and surface tension effects are not considered or limited by the contact model between the rigid body and the liquid provided by PhysX 5. The lack of these properties could reduce the realism of the simulations, potentially affecting the transferability of trained models to real-world scenarios. While our approach prioritizes computational feasibility and training efficiency, future work may explore other efficient soft body simulation techniques, such as the material point method (MPM) [49] to mitigate these limitations.
We have also considered the observation of historical images to capture the temporal information. Including historical images can potentially enhance the performance, such as by allowing the agent to infer liquid occluded in the current step by the Suction/Irrigator from previous non-occluded frames. However, considering the large rollout buffer size and the usage of demonstration data, a much larger amount of memory would be needed. In the current setting, the memory usage is already more than 30 gigabytes with IL for irrigation, making it impractical to stack the image observations with historical ones. To solve this issue, either a much larger memory should be used, or ML-Agents must be modified to allow streaming of the rollout data. Using historical frames at a lower resolution may also be beneficial, though further downscaling from the current $84\times 84$ resolution could result in significant information loss.
@@ VII-C Other Limitations and Future Work
In this work, no depth information is provided to the agent. Ideally, it may be possible to achieve a better performance when depth is taken into account. However, it is not included in this work as it may increase the sim-to-real gap, due to the additional noises and inaccuracies of the real-world depth images. Future work may either include the depth observation directly or train the agent with a stereo camera that captures two images side-by-side. Using a stereo camera can also provide additional information that a single camera might miss, in addition to depth, especially when tools occlude the liquid and tissue. However, in this work, significant tool occlusion is not encountered due to the tool’s relatively small radius and the side-view camera pose.
Other imitation learning approaches, such as diffusion policy [50] and action chunking with transformers [51, 52], can be further incorporated in future work. This may allow more in-depth investigation of pure IL approaches without RL, as well as novel approaches that combine the state-of-the-art IL methods with RL.
One limitation of the real-world experiments is that a hardcoded EE force constraint is imposed. If the force limit is reached when the agent attempts to move the EE toward a position, such as when the EE touches the tissue, the trial is terminated for safety. This happens only during suction, as close contact between the tool and tissue may be needed to suction the liquid. Among all the suction trials, including those within combined irrigation-suction, a total of 9 trials were terminated due to force violation. Although the agent is penalized when the EE is in contact with the tissue during training, force limit violations are still seen during some suction trials. Future work may employ a better training strategy, such as rewarding the agent for lifting up the EE when it is in contact with the tissue.
This work focuses primarily on obtaining agents that can perform the tasks autonomously under various physical configurations. While the evaluation focuses on task completion, other performance metrics are less considered, such as the execution time and the frequency of safety violations. Qualitative observations indicate that irrigation execution time remains stable across trials, whereas suction time varies depending on the complexity of the suction area. Further analysis of task failure cases and a comparison between agents’ performance and human operation may be necessary before considering surgical applications, which are beyond the scope of this research.
Further extension of this work to actual surgeries will also need additional development of realistic surgical scenes and re-training of the agents, as this work considers a relatively simple environment setup without actual surgical components, such as phantom tissue. With the simulation capability introduced in this work, transitioning to more realistic surgical scenes–including those beyond minimally invasive surgery–is straightforward, particularly when simulating body fluids. Additional validations are required as well to assess the performance in realistic environments, such as using cadavers.
@@ VIII Conclusion
In this work, we examined the problem of autonomous irrigation and suction in MIS using RL. To allow sim-to-real transfer, a new surgical robot learning simulation framework was built using Unity and PhysX 5, which enables the development of simulated learning environments for irrigation and suction. With DR and carefully designed reward functions, combined with CL and IL, two vision-based agents were trained to autonomously complete irrigation and suction. We found that the irrigation agent achieves a much better performance when a curriculum is used, and that IL helps improve the suction agent but not the irrigation one. We evaluated their performance in the real world and showed that they can achieve satisfactory results in most cases. Additionally, our simulation platform can be extended and used to build simulated training environments for other surgical tasks in the future.